OpenStudy (anonymous):

one quick question on polynomials ? (: Simplify : (4x − 6) + (5x + 1)

OpenStudy (whpalmer4):

Collect like terms. (4x - 6) + (5x + 1) = 4x - 6 + 5x + 1 = 4x + 5x + 1 - 6 = 9x - 5